ShadexX:

Simplify negative 2 and 1 over 6 − negative 7 and 1 over 3.

umm:

Hey there, @ShadexX, welcome to QuestionCove! Does it look like this?\[-2\frac{ 1 }{ 6 }-7\frac{ 1 }{ 3 }\]

ShadexX:

Thanks for the welcome! and yes, that's what it looks like.

umm:

No problem. Alright, let's start out with given, which is \[-2\frac{ 1 }{ 6 }-7\frac{ 1 }{ 3 }\] \[=-\frac{ 13 }{ 6 }-7\frac{ 1 }{ 3 }\] \[=\frac{ -13 }{ 6 }-7\frac{ 1 }{ 3 }\] \[=\frac{ -13 }{ 6 }-\frac{ 22 }{ 3 }\] \[=\frac{ -19 }{ 2 }\] \[=-9\frac{ 1 }{ 2 }\] Therefore your answer is:\[=-9\frac{ 1 }{ 2 }\]

ShadexX:

Wow, thanks! I really appreciate it! :D

umm:

Anytime, glad to be of assistance.